Solve the system by addition.
5x – 4y = –5
x – 2y = –7

Answer by stanbon(75887)

Solve the system by addition.
1st: 5x – 4y = –5
2nd: x – 2y = –7
-------------
Multiply 2nd by 2 to get:
3rd: 2x-4y=-14
Subtract 3rd from 1st to get:
3x=9
x=3
----------
Substitute into x-2y=-7 to solve for y:
3-2y=-7
2y = 10
y = 5
---------------
Solution: x=3 ; y=5
=============================
Cheers,
Stan H.